OpenStudy (anonymous):

(4x3 + 7x2 – 13x – 3) ÷ (x + 3) use synthetic division to solve this problem

OpenStudy (anonymous):

i got this so far 3) 4 7 -13 -3 12 -15 6 4 -5 2 -3

OpenStudy (anonymous):

Okk. but can you help me with this question.

OpenStudy (anonymous):

LOL YEA FAN ME

OpenStudy (anonymous):

i am afraid, my dear, you made the same mistake as last time

OpenStudy (anonymous):

remember when you divided by \(x-7\) and you had to make sure to put a \(7\) on the side, rather than \(-7\) ?

OpenStudy (anonymous):

in this case you have \(x+3\) so guess what you have to put on the side?

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okk thank you. One sec im going to try again.

OpenStudy (anonymous):

that is why you get to add instead of subtract when you do these that is the reason you want to put \(-3\) on the side, so you don't have to worry about subtracting like you do with long division

OpenStudy (anonymous):

Please see the attached file.

OpenStudy (anonymous):

-3) 4 7 -13 -3 -12 -15 -6 4 -5 -2 -3

OpenStudy (anonymous):

oohh thank you @adinarayan and @satellite73 . i see what ii did wrong . (: You guyrs were alot of help.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

No problem. Thanks for the compliment. :)

OpenStudy (anonymous):

Wait , sorry. so is this the answer ? x^3+4x^2-5x+2 R-9

OpenStudy (anonymous):

no i think there was a mistake in the answer given above

OpenStudy (anonymous):

:( . What was it ?

OpenStudy (anonymous):

lets re do it as you were doing it

OpenStudy (anonymous):

-3) 4 7 -13 -3 -12 -15 -6 4 -5 -2 -3

OpenStudy (anonymous):

4 7 -13 -3 -3 -12 15 -6 4 -5 2 -9

OpenStudy (anonymous):

your \(-15\) should be \(15\) and \(-13+15=2\)

OpenStudy (anonymous):

are you happy with my last line at the bottom?

OpenStudy (anonymous):

My gosh, I'm sorry. Did I give an incorrect answer? I honestly apologize, it won't happen again... :(

OpenStudy (anonymous):

if so, now that we have the last line as 4 -5 2 -9 we have to know how to read off the answer

OpenStudy (anonymous):

you have a polynomial of degree 3, divided by a polynomial of degree 1 leaving a polynomial of degree 2 the coefficients are there, so your answer is \[4x^2-5x+2\] with a remainder of \(-9\)

OpenStudy (anonymous):

okkk yes. i had got that @satellite73 but im not sure how to write it after i get that. and @adinarayan nooo its okk. i just was wondering how would i write the answer you agve me. x^3+4x^2-5x+2 R-9?

OpenStudy (anonymous):

you could write \[4x^2-5x+2-\frac{9}{x+3}\]

OpenStudy (anonymous):

@adinarayan i make mistakes all day, you will make more, i am sure of it

OpenStudy (anonymous):

i like to think of them as "typos"

OpenStudy (anonymous):

thank you. do you guys mind helping in another question?

OpenStudy (anonymous):

(3x3 – 4x – 1)÷ (x + 1)

OpenStudy (anonymous):

-1) 3 -4 -1 -3 1 3 -1 0 (3x^2-1)

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

No.\[3 x^3-4 x-1=(x+1) \left(3 x^2-3 x-1\right) \]

OpenStudy (anonymous):

another very very common mistake in this one you forgot to hole the place of the missing term

OpenStudy (anonymous):

sorry im confused.

OpenStudy (anonymous):

\[3x^3 – 4x – 1\] you need to put a \(0\) in the \(x^2\) place, as that term is missing

OpenStudy (anonymous):

when you list the coefficients, you have to keep the place of any missing power you are writing in descending order of the powers since there is no \(x^2\) term you have to hold the \(x^2\) place with a \(0\). more simply put, the coefficients are \[3, 0, -4, -1\]

OpenStudy (anonymous):

set up should look like 3 0 -4 -1 -1)

OpenStudy (anonymous):

it is a very common mistake, make sure to look out for it

OpenStudy (anonymous):

oooohhh. okkk . sooo -1) 3 0 -4 -1 -3 3 1 3 -3 -1 0

OpenStudy (anonymous):

got it !

OpenStudy (anonymous):

that is why the answer is what @robtobey wrote \[3x^2-3x-1\] and no remainder

OpenStudy (anonymous):

thankk youu . (:: . i just have three more queestions i need help on.

OpenStudy (anonymous):

post em

OpenStudy (anonymous):

make a new thread